# Preprint no. NJU-INP 105-25 Exclusive photoproduction of light and heavy vector mesons: thresholds to very high energies

### v1

@@ Abstract
A reaction model for $\gamma+p\to V+p$, $V=\rho^{0},\phi,J/\psi,\Upsilon$, which exposes the quark-antiquark content of the photon in making the transition $\gamma\to{\mathpzc q}\bar{\mathpzc q}+\mathbb{P}\to V$, where $\mathpzc q$} depends on $V$, and couples the intermediate ${\mathpzc q}\bar{\mathpzc q}$ system to the proton’s valence quarks via Pomeron ($\mathbb{P}$) exchange, is used to deliver a unified description of available data – both differential and total cross sections – from near threshold to very high energies, $W$, for all the $V$-mesons. For the $\Upsilon$, this means $10\lesssim W/{\rm GeV}\lesssim 2\,000$. Also provided are predictions for the power-law exponents that are empirically used to characterise the large-$W$ behaviour of the total cross sections and slope parameters characterising the near-threshold differential cross sections. Appealing to notions of vector meson dominance, the latter have been interpreted as vector-meson–proton scattering lengths. The body of results indicate that it is premature to link any $\gamma+p\to V+p$ data with, for instance, in-proton gluon distributions, the quantum chromodynamics trace anomaly, or pentaquark production. Further developments in reaction theory and higher precision data are required before the validity of any such links can be assessed.
@@ 1 Introduction
provides a clean environment to study diffractive scattering from the proton and thereby obtain information about the reaction mechanism and both hadrons involved in the process.
@@
Understanding the energy dependence of these cross sections has long challenged theory. Contrary to perturbative quantum chromodynamics (QCD) expectations, which suggest that total cross sections should fall as the center‑of‑mass energy-squared, $s=W^{2}$, increases, experiments show that such photoproduction cross sections actually rise with $s$ Baksay et al. (1978). This issue was analysed, e.g., in Ref. Foldy and Peierls (1963), with the conclusion that the mechanism responsible should correspond to exchange of something that has isospin zero and is even under charge conjugation. Interactions of this kind can be expressed in Regge phenomenology via Pomeron, $\mathbb{P}$, exchange, with a trajectory whose intercept $\alpha_{\mathbb{P}}(0)>1$. The QCD character of the Pomeron is widely discussed Donnachie et al. (2004); Gribov (2007). It is typically supposed to represent the collective exchange of a family of colourless, crossing-even states, with two-gluon exchange being the leading contributor Forshaw and Ross (2011). Connections with glueball physics are also discussed Meyer and Teper (2005); Boschi-Filho et al. (2006).
@@
Figure 1: Reaction model for $\gamma+p\to V+p$, referred to hereafter as $\mathbb{P}$-dyn. The quark-antiquark, ${\mathpzc q}\bar{\mathpzc q}$, component of the dressed-photon is probed by Pomeron exchange with the proton, producing an on-shell vector meson, $V$. Referring to Eq. (4): the green rectangle describes Pomeron exchange, $\Gamma^{\mathbb{P}}$, including its couplings to the proton and quark – see Sect. $2.$ Further, the solid black curve is $S_{q}$, the dressed $q$ quark propagator; the shaded lighter-blue circle is $\Gamma_{\mu}^{\gamma}$, the dressed-$\gamma q\bar{q}$ vertex; and the shaded red circle is $\Gamma_{\nu}$, the $V$-meson Bethe-Salpeter amplitude – see Sect. 3. Kinematics: $s=W^{2}=-(p_{2}^{2}+P^{2})$; $t=-(p_{2}-p_{1})^{2}$.
@@
With the high-energy behaviour of $\gamma+p\rightarrow V+p$ described by $\mathbb{P}$ exchange, it is natural to seek an explanation for the cross section at low-energies and determine whether a unified approach to the entire $s$-domain is possible. Both these issues were addressed in Ref. Pichowsky and Lee (1997), which explored the reaction mechanism depicted in Fig. 1 using phenomenological inputs for the required elements. This approach sees the ${\mathpzc q}\bar{\mathpzc q}$ component of the photon, where $\mathpzc q$ is chosen from the set $\{{\mathpzc q}=u/d,s,c,b\}$, depending on $V$, exposed by interaction with the proton via $\mathbb{P}$-exchange and then transformed into the on-shell vector meson when sufficient reaction energy is available, i.e., once the threshold energy is exceeded:
@@ 2.1 Dynamical photon to meson transition
where $N_{c}=3$; $e_{\rho^{0},\phi,J/\psi,\Upsilon}=e\{1/\surd 2,1/3,2/3,1/3\}$, the fine-structure constant $\alpha=e^{2}/[4\pi]$; the trace is over spinor indices; $k_{\pm}=k\pm P/2$; $\Gamma_{\nu}^{V}$ is the Bethe-Salpeter amplitude for an on-shell $V$ meson – see, e.g., Ref. Xu et al. (2021); $S_{\mathpzc q}$ is the dressed $q$ quark propagator; – see, e.g., Ref. (Roberts et al., 2021, Sect. 2.C); and $\Gamma_{\mu}^{\gamma}$ is the associated dressed photon-quark vertex. Since we are focusing on photoproduction, then, with complete generality, one may use the Ball-Chiu form for $\Gamma_{\mu}^{\gamma}$ Ball and Chiu (1980).
The remaining element in Eq. (4) is $\Gamma^{\mathbb{P}}_{\alpha}=\gamma_{\alpha}\beta_{\mathpzc q}$, the Pomeron-quark vertex. Here, $\beta_{\mathpzc q}$ is a constant strength factor and $\gamma_{\alpha}$ is a standard Dirac matrix. We have verified that replacing $\gamma_{\alpha}$ by some more sophisticated structure, such as the Ball-Chiu vertex Ball and Chiu (1980), for instance, has no material influence: any impact can be absorbed into a modest modification of $\beta_{\mathpzc q}$.
@@
$\frac{d\sigma}{d\Omega}=\frac{1}{4\pi^{2}}\frac{m_{p}}{4W}\frac{|\vec{P}|}{k_{W}}\sum_{\text{proton spin}}(|{\cal I}_{1}|^{2}+|{\cal I}_{2}|^{2})\,$, (7)
@@ 2.2 Pomeron in the reaction model
with $\alpha_{\mathbb{P}}(t)=\alpha_{0}+\alpha_{1}t$, $s_{0}=1/\alpha_{1}$, and $\alpha_{0,1}$ fitted to selected small-$|t|$, large-$W$ meson photoproduction data within the context of Regge phenomenology: $\alpha_{0}$ is fixed by the energy dependence of the total cross section and $\alpha_{1}$ by the $t$-slope of the differential cross section – see, e.g., Refs. Collins (1977); Irving and Worden (1977). These parameters depend on whether the final-state vector meson is light or heavy Lee et al. (2022).
Table 1: Quark current-masses and parameters characterising the CSM interaction kernels for the mesons considered herein. The quark current masses are quoted at a renormalisation scale $\zeta=\zeta_{19}=19\,$GeV, i.e., in the far ultraviolet so as to avoid truncation ambiguity. The $u,d$ values were chosen to reproduce $m_{\pi}$, $f_{\pi}$; the kaon mass was used to fix $m_{s}$; $m_{c}$ was chosen to yield the $\eta_{c}$ mass; and the $\eta_{b}$ mass was used to fix $m_{b}$ Qin et al. (2018). Using one-loop evolution, the listed light quark masses correspond to $\zeta=\zeta_{2}=2\,$GeV values of $m_{\ell}^{\zeta_{2}}=0.0048\,$GeV, $m_{s}^{\zeta_{2}}=0.117\,$GeV; and regarding the heavy quarks, the listed values equate to the following Euclidean constituent quark masses Hecht et al. (2001), $M_{c}=1.41\,$GeV, $M_{b}=4.34\,$GeV. All four masses correspond to those typically quoted in connection with the given flavour and are a fair match with contemporary inferences Navas et al. (2024). The same Pomeron trajectory is used for both $\rho^{0}$ and $\phi$ photoproduction Pichowsky and Lee (1997), and another for $J/\psi$ and $\Upsilon$ (Chekanov et al., 2002, ZEUS). Pomeron-quark couplings, $\beta_{\mathpzc q}$, for both reaction models considered herein – Sects. 2.1, 2.3 – are determined as described in Sect. 2.4.
@@ 2.3 VMD-like treatment of photon to meson transition
In the following, we will compare our predictions with those obtained using a simplified framework that replaces the calculated $\gamma\to{\mathpzc q}\bar{\mathpzc q}+\mathbb{P}\to V$ transition matrix element by a phenomenologically constrained model. Sketched in Fig. 2 and identified as the $\mathbb{P}\,$-am model in Ref. Tang et al. (2024), this approach is obtained by beginning with the Ref. Sakinah et al. (2024) model; setting $v_{cN}=0$, so the quark-loop integration, based on a constituent quark model wavefunction, and VNV-N final state interactions (FSIs) do not contribute to the amplitude; and updating the Pomeron trajectories and Pomeron-quark couplings.
The associated matrix element for diffractive vector meson photoproduction is
@@
with the tensor structure of the $\mathbb{P}$-$V$ vertex being Titov et al. (1998)
@@ 3 Photon to vector meson transition amplitude in CSMs
$k^{2}T_{\mu\nu}(k)=k^{2}\delta_{\mu\nu}-k_{\mu}k_{\nu}$, $y=k^{2}$. The tensor structure corresponds to Landau gauge, used because it is both a fixed point of the renormalisation group and that gauge for which corrections to RL truncation are least noticeable Bashir et al. (2009). In Eq. (3), $r,s,t,u$ represent colour, spinor, flavour matrix indices (as necessary).
A realistic form of $\tilde{\mathpzc G}(y)$ is explained in Refs. Qin et al. (2011); Binosi et al. (2015):
@@ 4 Near-threshold slope parameter / meson-proton scattering lengths
$\frac{d\sigma^{\gamma p\to Vp}}{d\Omega}\bigg|_{\text{th}}$ $=$ $\frac{|\vec{P}|}{|\vec{q}|}\frac{1}{64\pi}|T^{\gamma p\to Vp}|^{2},$ (18)
$=$ $\frac{|\vec{P}|}{|\vec{q}|}\frac{\pi\alpha}{g_{V}^{2}}\frac{d\sigma^{Vp\to Vp}}{d\Omega}\bigg|_{\text{th}}$
$=$ $\frac{|\vec{P}|}{|\vec{q}|}\frac{\pi\alpha}{g_{V}^{2}}|\alpha_{Vp}|^{2},$
where $\alpha$ is the fine-structure constant, $\vec{q}$ and $\vec{P}$ are the initial and final momenta of the photon and vector meson, respectively – Eq.(5), and $\alpha_{Vp}$ is the model meson-proton scattering length. The VMD ingredient is the coupling $g_{V}$, which is defined by the vector meson’s electromagnetic decay width $\Gamma_{V\to e^{+}e^{-}}$:
@@
Our prediction for this differential cross section is given in Eq. (7). That is the formula we use to evaluate all results. However, by equating our predictions with Eq. (18) one can arrive at model-specific VMD values for the meson-proton scattering length. These can be used as single parameters with which to evaluate the veracity of existing such model dependent inferences of meson-proton scattering lengths Strakovsky et al. (2020a, b, 2021, 2025). Furthermore, obtained in this way, one need not interpret $|\alpha_{Vp}|$ as the VpV-p scattering length. Instead, it can be viewed as a near-threshold slope parameter. Then, the value has objective utility as a comparison between theory and experiment.

### v2

# Preprint no. NJU-INP 105/25 Exclusive photoproduction of light and heavy vector mesons: thresholds to very high energies

@@ Abstract
A reaction model for $\gamma+p\to V+p$, $V=\rho^{0},\phi,J/\psi,\Upsilon$, which exposes the quark-antiquark content of the photon in making the transition $\gamma\to{\mathpzc q}\bar{\mathpzc q}+\mathbb{P}\to V$, where $\mathpzc q$} depends on $V$, and couples the intermediate ${\mathpzc q}\bar{\mathpzc q}$ system to the proton’s valence quarks via Pomeron ($\mathbb{P}$) exchange, is used to deliver a unified description of available data – both differential and total cross sections – from near threshold to very high energies, $W$, for all the $V$-mesons. For the $\Upsilon$, this means $10\lesssim W/{\rm GeV}\lesssim 2\,000$. Also provided are predictions for the power-law exponents that are empirically used to characterise the large-$W$ behaviour of the total cross sections and slope parameters characterising the near-threshold differential cross sections. Appealing to notions of vector meson dominance, the latter have been interpreted as vector-meson–proton scattering lengths. The body of results indicates that it is premature to link any $\gamma+p\to V+p$ data with, for instance, in-proton gluon distributions, the quantum chromodynamics trace anomaly, or pentaquark production. Further developments in reaction theory and higher precision data are required before the validity of any such links can be assessed.
@@ 1 Introduction
provides a clean environment to study diffractive scattering from the proton and thereby obtain information about the reaction mechanism and both hadrons involved in the process. (Herein, we consider only ground-state vector-mesons.)
@@
Understanding the energy dependence of these cross sections has long challenged theory. Contrary to perturbative quantum chromodynamics (QCD) expectations, which suggest that total cross sections should fall as the center‑of‑mass energy-squared, $s=W^{2}$, increases, experiments show that such photoproduction cross sections actually rise with $s$ Baksay et al. (1978). This issue was analysed, e.g., in Ref. Foldy and Peierls (1963), with the conclusion that the mechanism responsible should correspond to exchange of something that has isospin zero and is even under charge conjugation. Interactions of this kind can be expressed in Regge phenomenology via Pomeron, $\mathbb{P}$, exchange, with a trajectory whose intercept $\alpha_{\mathbb{P}}(0)>1$.
The QCD character of the Pomeron is widely discussed Donnachie et al. (2004); Gribov (2007). It is typically supposed to represent the collective exchange of a family of colourless, crossing-even states, with two-gluon exchange being the leading contributor Forshaw and Ross (2011). Some phenomenological approaches distinguish between a hard (not necessarily perturbative) and soft Pomeron Donnachie and Landshoff (1998); and connections with glueball physics are also discussed Meyer and Teper (2005); Boschi-Filho et al. (2006).
@@
Figure 1: Reaction model for $\gamma+p\to V+p$, referred to hereafter as $\mathbb{P}$-dyn. The quark-antiquark, ${\mathpzc q}\bar{\mathpzc q}$, component of the dressed-photon is probed by Pomeron exchange with the proton, producing an on-shell vector meson, $V$. Referring to Eq. (4): the green rectangle describes Pomeron exchange, $\Gamma^{\mathbb{P}}$, including its couplings to the proton and quark – see Sect. $2.$ Further, the solid black curve is $S_{q}$, the dressed $q$ quark propagator; the shaded lighter-blue circle is $\Gamma_{\mu}^{\gamma}$, the dressed-$\gamma q\bar{q}$ vertex; and the shaded red circle is $\Gamma_{\nu}$, the $V$-meson Bethe-Salpeter amplitude – see Sect. 3. Kinematics: $s=W^{2}=-(p_{2}+P)^{2}$; $t=-(p_{2}-p_{1})^{2}$.
Twenty and more years ago, a variety of such models were proposed as motivations for the measurement of $\gamma+p\rightarrow V+p$ photoproduction reactions; see, e.g., Refs. Nemchik et al. (1994); Cudell and Royen (1997); Frankfurt et al. (1999); Brodsky et al. (2001); Forshaw et al. (2004). A number were built upon the assumption of factorisation for hard exclusive processes, e.g., Nemchik et al. (1994); Frankfurt et al. (1999); Brodsky et al. (2001), and different studies canvassed the possibilities of distinct microscopic interaction mechanisms. Notwithstanding that, all such models attempted to draw a connection between $V$-meson photoproduction reactions and features of the target and produced meson, like in-target gluon structure functions, with such links drawn using Ansätze for the light-front wave functions of the bound states involved. Infrared freezing of the QCD running coupling was also sometimes assumed in the interaction models Nemchik et al. (1994); Cudell and Royen (1997), although with saturation values chosen which are significantly less than that associated with QCD’s process-independent effective charge Cui et al. (2020); Deur et al. (2024); Brodsky et al. (2024). Owing to their formulation assumptions, the models were only considered plausible at energies far above threshold. As will become clear, our framework represents a significant departure from such approaches. For instance, inter alia: no call is made to factorisation for hard exclusive processes; the required bound-state wave functions and constituent propagators are consistently calculated using an interaction that derives from modern analyses of QCD’s gauge sector; and predictions are simultaneously made for all $V$-meson final states, at energies that range, in each case, from threshold to very far above.
@@
Continuing with this last point, having noted that the high-energy behaviour of $\gamma+p\rightarrow V+p$ is described by $\mathbb{P}$ exchange, it is natural, as herein, to seek an explanation for the cross section at low-energies and determine whether a unified approach to the entire $s$-domain is possible. Both these issues were addressed in Ref. Pichowsky and Lee (1997), which explored the reaction mechanism depicted in Fig. 1 using phenomenological inputs for the required elements. This approach sees the ${\mathpzc q}\bar{\mathpzc q}$ component of the photon, where $\mathpzc q$ is chosen from the set $\{{\mathpzc q}=u/d,s,c,b\}$, depending on $V$, exposed by interaction with the proton via $\mathbb{P}$-exchange and then transformed into the on-shell vector meson when sufficient reaction energy is available, i.e., once the threshold energy is exceeded:
@@ 2.1 Dynamical photon to meson transition
where $N_{c}=3$; $e_{\rho^{0},\phi,J/\psi,\Upsilon}=e\{1/\surd 2,1/3,2/3,1/3\}$, the fine-structure constant $\alpha=e^{2}/[4\pi]$; the trace is over spinor indices; $k_{\pm}=k\pm P/2$; $\Gamma_{\nu}^{V}$ is the Bethe-Salpeter amplitude for an on-shell $V$ meson – see, e.g., Ref. Xu et al. (2021); $S_{\mathpzc q}$ is the dressed $q$ quark propagator; – see, e.g., Ref. (Roberts et al., 2021, Sect. 2.C); and $\Gamma_{\mu}^{\gamma}$ is the associated dressed photon-quark vertex. Since we are focusing on photoproduction, then, with complete generality, one may use the Ball-Chiu form for $\Gamma_{\mu}^{\gamma}$ Ball and Chiu (1980):
$i\Gamma_{\mu}^{\rm BC}$ $(k_{1},k_{2})=\Sigma_{A}(k_{1}^{2},k_{2}^{2}))i\gamma_{\mu}+(k_{1}+k_{2})_{\mu}$
$\times[\Delta_{A}(k_{1}^{2},k_{2}^{2})i\gamma\cdot(k_{1}+k_{2})+\Delta_{B}(k_{1}^{2},k_{2}^{2})]\,$, (5)
where $\Sigma_{F}(k_{1}^{2},k_{2}^{2})=[F(k_{1})^{2}+F(k_{2}^{2})]/2$, $\Delta_{F}(k_{1}^{2},k_{2}^{2}))=[F(k_{1})^{2}-F(k_{2}^{2})]/[k_{1}^{2}-k_{2}^{2}]$, $F=A,B$. This Ansatz is complete for photoproduction because, with $Q^{2}=0$, it delivers the Ward identity result: $i\Gamma_{\mu}^{\rm BC}(k_{-},k_{-})=\partial_{\mu}S(k_{-})$.
The remaining element in Eq. (4) is $\Gamma^{\mathbb{P}}_{\alpha}$, the Pomeron-quark vertex. Following Refs. Donnachie and Landshoff (1984); Pichowsky and Lee (1997) and for the reasons detailed therein, we employ $\Gamma^{\mathbb{P}}_{\alpha}=\gamma_{\alpha}\beta_{\mathpzc q}$, where $\beta_{\mathpzc q}$ is a constant strength factor and $\gamma_{\alpha}$ is a standard Dirac matrix.
[…]
@@
[…]
@@ 2.2 Pomeron in the reaction model
with $\alpha_{\mathbb{P}}(t)=\alpha_{0}+\alpha_{1}t$, $s_{0}=1/\alpha_{1}$, and $\alpha_{0,1}$ fitted to selected small-$|t|$, large-$W$ meson photoproduction data within the context of Regge phenomenology: $\alpha_{0}$ is fixed by the energy dependence of the total cross section and $\alpha_{1}$ by the $t$-slope of the differential cross section – see, e.g., Refs. Collins (1977); Irving and Worden (1977).
As observed, e.g., in Refs. Donnachie and Landshoff (1998); Pichowsky and Lee (1997); Lee et al. (2022), data suggests that in comparing photoproduction reactions involving light or heavy vector-meson final states, different values of the pair $(\alpha_{0},\alpha_{1})$ are required; see, also, below. This observation encouraged some practitioners to suggest the existence of two Pomerons Donnachie and Landshoff (1998), one hard and one soft, with both emerging as the result of some (unknown) nonperturbative dynamics. Given the current QCD understanding of Pomeron phenomenology and structure, we elect not to express an opinion on this matter. Instead, in the reaction model advocated herein, there is a single (soft) exchanged object. The object fits within Pomeron-based schemes because, amongst other things, it provides a nonlocal color-singlet exchange that produces cross sections which rise with energy. This Pomeron represents essentially nonperturbative aspects of reactions involving hadrons; and the proton + Pomeron + quark interaction possesses an energy dependence that differs between light-quark meson and heavy-quark meson final states. It is, perhaps, worth noting that scale-dependent Pomeron trajectory parameters may be natural in some approaches, e.g., models based on AdS/CFT duality Dosch and Ferreira (2015).
Table 1: All parameters used herein. Quark current-masses and parameters characterising the CSM interaction kernels for the mesons considered herein. The quark current masses are quoted at a renormalisation scale $\zeta=\zeta_{19}=19\,$GeV, i.e., in the far ultraviolet so as to avoid truncation ambiguity. The $u,d$ values were chosen to reproduce $m_{\pi}$, $f_{\pi}$; the kaon mass was used to fix $m_{s}$; $m_{c}$ was chosen to yield the $\eta_{c}$ mass; and the $\eta_{b}$ mass was used to fix $m_{b}$ Qin et al. (2018). Using one-loop evolution, the listed light quark masses correspond to $\zeta=\zeta_{2}=2\,$GeV values of $m_{\ell}^{\zeta_{2}}=0.0048\,$GeV, $m_{s}^{\zeta_{2}}=0.117\,$GeV; and regarding the heavy quarks, the listed values equate to the following Euclidean constituent quark masses Hecht et al. (2001), $M_{c}=1.41\,$GeV, $M_{b}=4.34\,$GeV. All four masses correspond to those typically quoted in connection with the given flavour and are a fair match with contemporary inferences Navas et al. (2024). The same Pomeron trajectory is used for both $\rho^{0}$ and $\phi$ photoproduction (Pichowsky and Lee, 1997, Sec. III.A), and another for $J/\psi$ and $\Upsilon$ (Chekanov et al., 2002, ZEUS). Pomeron-quark couplings, $\beta_{\mathpzc q}$, for both reaction models considered herein – Sects. 2.1, 2.3 – are determined as described in Sect. 2.4.
@@ 2.3 VMD-like treatment of photon to meson transition
In the following, we will compare our predictions with those obtained using a simplified framework that replaces the calculated $\gamma\to{\mathpzc q}\bar{\mathpzc q}+\mathbb{P}\to V$ transition matrix element by a phenomenologically constrained model. Sketched in Fig. 2 and identified as the $\mathbb{P}\,$-am model in Ref. Tang et al. (2024), this approach is obtained by beginning with the Ref. Sakinah et al. (2024) model; setting $v_{cN}=0$, so the quark-loop integration, based on a constituent quark model wavefunction, and VNV-N final-state interactions (FSIs) do not contribute to the amplitude; and updating the Pomeron trajectories and Pomeron-quark couplings.
The associated matrix element for diffractive vector meson photoproduction is (Sakinah et al., 2024, Eqs. (41-43)):
@@
with the tensor structure of the $\mathbb{P}$-$V$ vertex being a minimal Ansatz that ensures physical constraints (Titov et al., 1998, Sec. III.A), viz.
@@ 3 Photon to vector meson transition amplitude in CSMs
$k^{2}T_{\mu\nu}(k)=k^{2}\delta_{\mu\nu}-k_{\mu}k_{\nu}$, $y=k^{2}$. The tensor structure corresponds to Landau gauge, used because it is both a fixed point of the renormalisation group and that gauge for which corrections to RL truncation are least noticeable Bashir et al. (2009). In Eq. (3), $r,s,t,u$ represent colour, spinor, flavour matrix indices (as necessary). A realistic form of $\tilde{\mathpzc G}(y)$ is explained in Refs. Qin et al. (2011); Binosi et al. (2015):
@@
At this point, with the quark current masses and parameter values in Table 1 left-block, the necessary gap and Bethe-Salpeter equations can be solved using standard algorithms Maris and Roberts (1997); Maris and Tandy (2006); Krassnigg (2008) to obtain each element in Eq. (4) and therewith $t_{\mu\alpha\nu}^{V}(q,P)$ itself. In solving all relevant Schwinger function equations, we use a mass-independent (chiral-limit) momentum-subtraction renormalisation scheme Chang et al. (2009), with renormalisation scale $\zeta=19\,$GeV. At this scale, the quark wave function renormalisation constants are practically unity.
@@ 4 Near-threshold slope parameter / meson-proton scattering lengths
$\frac{d\sigma^{\gamma p\to Vp}}{d\Omega}\bigg|_{\text{th}}$ $=$ $\frac{|\vec{P}|}{k_{W}}\frac{1}{64\pi}|T^{\gamma p\to Vp}|^{2},$ (19)
$=$ $\frac{|\vec{P}|}{k_{W}}\frac{\pi\alpha}{g_{V}^{2}}\frac{d\sigma^{Vp\to Vp}}{d\Omega}\bigg|_{\text{th}}$
$=$ $\frac{|\vec{P}|}{k_{W}}\frac{\pi\alpha}{g_{V}^{2}}|\alpha_{Vp}|^{2},$
where $\alpha$ is the fine-structure constant, $\vec{P}$ is the final momentum of the vector meson – Eq.(6), and $\alpha_{Vp}$ is the model meson-proton scattering length. The VMD ingredient is the coupling $g_{V}$, which is defined by the vector meson’s electromagnetic decay width $\Gamma_{V\to e^{+}e^{-}}$:
@@
Suppose now that $V$-meson production near threshold is purely an $S$-wave process, then $\sigma_{\rm th}=4\pi d\sigma/d\Omega|_{\rm th}$. Then if one has precise total cross section data in hand, very close to threshold, it is possible to fit that data using a polynomial in $|\vec{P}|$, viz. write Strakovsky et al. (2020a)
$\sigma_{\rm th}={\mathpzc a}_{1}|\vec{P}|+{\mathpzc a}_{3}|\vec{P}|^{3}+{\mathpzc a}_{5}|\vec{P}|^{5},$ (21)
whereafter one has
$|\alpha_{Vp}|^{2}={\mathpzc a}_{1}\frac{g_{V}^{2}k_{W}}{4\pi^{2}\alpha}\,.$ (22)
Our prediction for the differential cross section is given in Eq. (8). That is the formula we use to evaluate all results. However, by equating Eq. (8) with Eq. (19) one can arrive at model-specific VMD values for the meson-proton scattering length:
$|\alpha_{Vp}|^{2}=\frac{1}{16\pi^{2}}\frac{m_{p}}{W}\frac{g_{V}^{2}}{\pi\alpha}\,\tfrac{1}{2}\!\!\!\sum_{\text{proton spin}}\!\!\!\left.(|{\cal I}_{1}|^{2}+|{\cal I}_{2}|^{2})\right|_{\rm th}\,.$ (23)
These can be used as single parameters with which to evaluate the veracity of existing such model dependent inferences of meson-proton scattering lengths Strakovsky et al. (2020a, b, 2021, 2025). Furthermore, obtained in this way, one need not interpret $|\alpha_{Vp}|$ as the VpV-p scattering length. Instead, it can be viewed as a near-threshold slope parameter. Then, the value has objective utility as a comparison between theory and experiment. We stress that one must use data very close to threshold if the fitted value of $\mathpzc a_{1}$ in Eq. (21) is to deliver a reliable empirical result for the slope parameter in Eq. (23).